Weekly Science Report 3: LMG02-05 Southern Ocean GLOBEC

Cruise days 12 to 18 August inclusive


BG-232-0 Costa/Burns/Crocker – Foraging Ecology of Crabeater seals

BG-234-0 Fraser – Winter Foraging Ecology of Adélie Penguins

BG-235-0 Fritsen – Sea Ice Microbial Communities

OG-241-0 Smith/Martinson/Perovich   Optical Environment of the Western Antarctic Peninsula Region

BG-244-0 Quetin/Ross – Winter Ecology of Larval Krill: Quantifying their interaction with the Pack Ice habitat


Cruise overview to date:


We began the week by finishing our convoy to the southern portion of the GLOBEC grid with the NBP and selecting a seemingly stable floe as Process Site #1 (informally dubbed Ice Station Sparky).  The Ice and Krill teams conducted initial surveys before deciding this was the place to stay.  We planned to conduct morning and evening CTD casts from the fantail, but rather quickly ran into problems - the ice rapidly closed off the open water we expected to have, and the grease and brash ice clogged the CTD pumps.  We tried several things, and finally settled on the solutions of keeping the CTD warm in the Baltic room between casts, turning on the pumps after lowering the package well beneath the ice cover, and using the ship's yokohama boat fender as a “plug” to reduce ice formation between casts.  Eventually even these methods didn't work due to relative motion of the ship and the ice and we began to use a SeaCat and XBTs through a 10" auger hole instead.  In the meantime, daily operations included scuba diving, ice coring, and seal captures.  The Seal team enjoyed unexpected success and was able to complete the deployment of their 11 satellite tags this week.  The Ice team deployed the first of their three buoy/sensor suites.  Wednesday the 14th brought cold temperatures and high winds, preventing most outdoor activities.  However, the scientists remained busy in the labs analyzing samples and processing data.  Somewhat surprisingly, our selected floe held together for the most part, though some active ridging consumed most of the original ice thickness transect and the primary dive hole.  At week's end we completed operations at this site, met up with the NBP, and began our convoy towards Process Site #2 offshore of the mouth of Marguerite Bay.


Daily Activities:


12 August 2002

The LMG has firmly established Process Site #1, informally dubbed “Sparky” in honor of our MT whose birthday was yesterday.  This morning two minke whales were observed spy hopping and several seals have briefly hauled out in our vicinity.  The first scuba dive of the day was cut short by a visit from a leopard seal; no injuries were sustained.  The ice team has been quite successful in completing an initial snow/ice transect and they are planning to extend their lines.  We conducted a CTD only (without rosette bottles) cast both last evening and this morning.  We plan to complete two casts per day, 12 hours apart.  Only the penguin team has been without luck, but they are assisting other groups with their activities.


13 August 2002

The LMG continues to occupy Process Station #1.  Daily operations include scuba diving, ice coring, seal work, and twice daily CTD casts to 300 meters.  The CRREL personnel deployed one of their three ice drifter buoys in the afternoon, which will remain in place after we leave, and relay information on ice drift, meteorological conditions, and temperature profiles through the ice. The seal team was very busy with numerous seals both hauled out on the ice and also playfully swimming around. The diving team had help on their dives today as the seals were using the same holes in the ice as the divers. Some of the seals would come out on top of the ice to assist the dive tenders.


14 August 2002

The LMG continues to occupy Process Station #1.  Yesterday ended successfully with all groups completing activities and collecting data.  Our first ice buoy was also deployed.  Today all activities have been aboard the ship as the temperature has dropped and the wind speed increased.  We are now drifting with the pack ice to the NNE at about 1.0 knot but there has been relatively little breakup of the ice. Today was the coldest and windiest day of the cruise so far, with temperatures hovering around -18ºC and winds gusting up to 40 knots.  The high winds pushed the ship and floe North at a speed of 1.5 knots during most of the day, keeping the divers out of the water and the seals under the ice.  One lone, brave Emperor penguin was sited speeding on his belly across the ice floe.


As evening (and darkness) fell the winds died down to 15 knots, and the Ice Team decided to brave the elements to go out onto the ice.  With the help of the ship lights they cored a second-year ice floe that was 2.5 m thick with more than 1.3m of snow and slush on top.  All in all it was another successful day for scientists aboard the L. M. Gould.


15 August 2002

The LMG continues to occupy Process Station #1.  Our drift has slowed from as much as 1.1 knot yesterday afternoon to 0 today as the winds dropped and the barometer stabilized.  All teams except those looking for penguins are able to collect samples and data today, and the seal team deployed the 10th of their 11 total tags this afternoon.  This has been a very productive location.


16 August 2002

The LMG continues to occupy Process Station #1.  Our drift speed has decreased to near zero and the weather has held steady, allowing all groups to access the ice and collect data except for the penguin team.  No birds besides a few snow petrels have been sighted recently.  However, the seal team deployed their 11th and final satellite tag this morning. As there were many seals in the area today the seal team carried out their first stomach flushing and found nothing in the stomach of a 12th crabeater seal. The krill team continues to conduct scuba dives, and the ice team is getting lots of cores to sample.


17 August 2002  

The LMG is completing our final day at Process Station #1 before moving on.  Normal operations include scuba diving and ice coring.  We have been shooting an XBT in the morning and another in the evening in lieu of a regular CTD cast.  Ice conditions and our concern about breaking up our working floe have prevented us from keeping a CTD hole open.  Although this has been a productive site, very few krill have been found either in the water, under the ice, or in stomach samples collected from predators.  This is curious, given the relative abundance of seals we've been seeing.


18 August 2002

We left ice station Sparky late last night and began our transit to our next process station.  We found the ice conditions favorable for travel over most of the route.  Once penguin was sited and an attempt was made to put the penguin team on the ice, however, the bird disappeared in the ice fog before we could get the penguin capture team on the ice.


Individual group reports: 


BG-232 – Seal Foraging Ecology- (Costa, Shaffer, Barnes, McDonald, and Kuhn)

This was a banner week for the seal team!  At ice station Sparky we completed the deployment of our 11th Satellite linked time depth recorder!  On 12 August we captured and tagged a 247 kg male, on 13 August a 301.5 kg male, on 15 August a 269 kg male and on 16 August we caught two 224 kg males, one of them received a satellite tag and the other was used for stomach sampling.  We now have 5 satellite linked depth recorders on 5 female and 5 male crabeater seals and one leopard seal.  All tags are working and we are receiving data on their movement patterns and dive behavior. We have completed morphometric, body composition and blood volume data on all of animals captured.  We plan to continue to capture seals for stomach sampling as well as to collect additional body composition data. 


BG-234 – Penguin Foraging Ecology (Heidi Geisz , Brett Pickering)

The first ice station did not prove to be productive for Adélie Penguin work; in fact none were seen during the week.  Several Snow Petrels, two Giant Petrels and a single Emperor Penguin were noted while on site.    The penguin team spent the week assisting others in the field, especially BG-232 with their seal work.  We also helped BG-235 core ice and dive tended for BG-244.  While not in the field we also tested transmitters, maintained equipment and plotted positions for penguins that were tagged earlier during this cruise.  


BG-235 - Sea Ice Microbial Communities (Fritsen, Hartsough, Adkins, Cunningham & Blees)

Work continued this week on ice station Sparky.  Coring activities produced ice with chlorophyll a concentrations ranging between 0.8 to 120 micrograms per liter.  Dark bands of algae found in several of the cores were located at depths ranging from 50 to 80 cm in the ice column but were not generally at the bottom of the ice cover.  Rather they were located within the interior of the ice.   Photosynthesis-irradiance assays indicated that these internal assemblages (which appeared to be dominated by pennate diatoms) were photoadapted to irradiances of ca. 50 micro mol photons per meter squared and capable of in situ growth.  Experimentation continued on algal growth potential, bacterial production and community analysis on ice and brine


OG-241 Optical Environment of the Western Antarctic Peninsula Region (Elder and Claffey)

While at process station “Sparky” we studied the different ice types in the area of the ship in conjunction with BG-235 and BG-244.  We performed ice transects for a total of 94 meters, a 64 m transect with a 30 meter “dog-leg” in the middle. Every meter along the transect, total snow depth, depth of liquid surface water (Slush), depth of the wicked layer (layer of snow that has wicked up moisture), snow/ice interface temperature, ice thickness and freeboard (the comparison of the top of the ice with sea level) were measured.  These transects were performed on the first days of the ice station. The first 10 meters of the transect were re-measured on the last day of the station. We found that the surface slush had frozen into ice but there was no measurable gain or loss in ice thickness along the bottom of the ice floe.  The ice around station “Sparky” was very active. Many times throughout the week, the floe cracked and rafted, thus the only section of the transect line that had not been altered due to ice motion was this first 10 meters.

Ice cores were taken at various locations. We measured temperature and salinity and d18O every 5 to 10 cm along the ice core.  Sections of core were also brought back to the ship for ice crystal structure measurements. Along with each set of these cores a snow pit was performed with measurements of temperature, salinity, grain size, type, and density.


Ice optics were performed on 2 different ice floes, 60cm and 35cm thick ice. These measurements determine the level of light that makes it through the snow and ice.  We measure incoming solar radiation, albedo (the reflection off the snow surface), and the transmission of light through the snow and ice.   Then we shovel the entire area and measure albedo and transmission of just the ice. This allows us to better separate the snow and ice portions of the light budget. A set of ice cores and snow pits were taken at each of these sites plus an additional core that will be sent back to the lab for di-electric measurements.


A drifting ice station was deployed.  Data from this drifter is sent back to our laboratory through the ARGOS satellite system.  On this drifter, we measure, air temperature, barometric pressure, GPS positions, snow depth, ice growth/decay on the under side of the ice, and ice buoyancy.   Also attached are 44 temperature sensors spaced at 5cm intervals along the vertical rod, positioned such that we have air, snow, ice, and ocean temperatures.  These data are updated every 2 hours.


While on station we maintained the passive microwave system to record a time series of measurements on the same piece of ice with changes in temperature. While underway, we video the ice along the ship track to help in the analysis of these data.


BG-244 - Winter Ecology of Larval Krill: Quantifying their interaction with the Pack Ice habitat (Quetin, Hessell, Willis, Boch, Oakes, Johnston and Dovel)

The past week was spent at the first Process Station (Ice Station Sparky).  This station was very dynamic the first several days so we did not install permanent underwater transects.  For a while it seemed we were cutting a new dive hole each morning.  Mid-week the wind shifted and temperatures dropped with the result that the ice was much less prone to cracking and over-rafting.  We lost one day of diving due to weather.  We had another exciting day diving with crabeater seals that were comfortable sharing our dive holes.  Luckily we cut the dive hole large enough for a diver and a seal!  This seal-human interaction proved beneficial to both parties, since the seals would keep our holes ice-free at night, and our awkward attempts to move efficiently through the water would provide them a few hours of entertainment during the day.


On 8/13 we observed an aggregation of larval krill in the water column below the sea ice.  The size of the aggregation (very approximately) was 10x15 m and 3-4 m deep, starting 2-3 m below the pack ice.  Based on our experience from past winter cruises and the GLOBEC cruise last winter, this is unusual behavior for krill larvae. Usually the larvae are actively feeding on or in closer proximity to the under ice surfaces.   No other krill were seen during the week, which is also very different from our experience last season when we observed larvae associated with the ice every day we dove.  We sampled larvae for a growth experiment, chn and length and stage frequency.  The larvae sampled ranged from F3 to juvenile stages, quite a wide range for this time of year.


Sea ice at the ice station was over-rafted to 12 m with mostly new over-rafting.  Visually, though most of the ice was white (indicating low concentrations of ice algae), there were some floes that looked yellow/brown, containing high concentrations of micro plankton communities.  This observation is also different from last winter when the ice was more uniformly contained low concentrations of ice algae (more uniformly white).  However, we were unable to dive this far south or this far offshore last season so comparisons with last year are difficult.  The next ice station located near Station 41 should provide a better comparison with last with season.  Divers collected approximately 40 samples from the over-rafted and floe ice surfaces for hplc, chn, chl and microscopic analyses.  In addition, several ice cores collected in coordination with Chris Fritsen's project will be analyzed in a similar manner.


The week went well except for the shifting ice early in the week and the lack of krill larvae.  We have a great group on board, from the bridge to the engine room.  Meals are digested with laughter, and the enthusiasm for the science is contagious as shown by some of the crew sitting at our dive hole and the Captain checking on the seal operation.


Submitted by:

Daniel Costa, Chief Scientist and

Karl Newyear, Marine Projects Coordinator

LMG 02-05